OpenStudy (anonymous):

(1 pt) Find the side length of the cube with the given volume. V= o.125 CM3

OpenStudy (wolfe8):

Use the formula of V=l^3 Meaning the volume of a cube is the cube of its length.

OpenStudy (anonymous):

cube volume = side^3 side = volume^1/3

OpenStudy (anonymous):

side = \[\sqrt[3]{0.125}\] = 0.5

OpenStudy (anonymous):

thanks joelmathews1234 :D